# Multilingual State Space Models for Structured Question Answering in Indic Languages

### v1

@@ 3.4 Prompting Strategies for Indic Language based SSMs
SSMs like Mamba and Mamba-2 require careful prompt engineering to maximize the question-answering capabilities in low-resource languages such as Hindi and Marathi. Also because of their lexicals , the tokens are not of constant length. Our approach combines the architectural strengths of SSMs with linguistic consideration for Devanagari script based languages.
@@
For Few-shot prompting Ye & Durrett (2022), The model maintains separate state trajectories for example patterns through its recurrent SSM block.
Figure 3: This example shows the one shot prompt in Hindi where we ask Question and give a scenario and then ask the question based on it.
We provide examples of prompts with both positive and negative samples in Appendix  Figure 8, showcasing their use in few-shot prompting, while examples for one-shot prompting are illustrated in  Figure 3. Our Question-Answering pipeline is based on the few-shot prompting framework  (Reynolds & McDonell, 2021). This approach demonstrates how leveraging paradigms from different prompting techniques leads to improved results during fine-tuning, as evidenced in the subsequent  Section 4.
We tend to select the best answer in P where $P\in K;K\geq 0$, are the number of samples generated for that query. These queries are converted into positive and negative sample. These selection of positive and negative samples are done using infomax algorithm Veličković et al. (2018) to retrieve the best answer out of $P$ samples, which makes the system stochastic and more deterministic.
@@ 3.5 Evaluation Metrics
To comprehensively evaluate model performance in question-answering tasks, we employed multiple metrics capturing different aspects of accuracy and relevance. EM provides a strict precision measure by evaluating the percentage of predictions that exactly match the ground truth, making it ideal for extractive QA tasks (Maalouly, 2022). F1 Score balances precision and recall, rewarding partial matches to better reflect token-level similarity, which is especially useful for paraphrased responses (Powers, 2020). To assess deeper semantic relationships, BERTScore (Zhang et al., 2020) leverages contextual embeddings from a pre-trained model, capturing meaning beyond exact word matches. ROUGE evaluates n-gram recall and overlap, ensuring that key phrases from the reference answer are retained even if the phrasing differs (Lin, 2004). Lastly, BLEU quantifies n-gram precision while incorporating a brevity penalty to prevent overly short predictions, making it particularly effective for assessing fluency and word-level accuracy in extractive QA settings (Papineni et al., 2002). These metrics collectively provide a robust assessment of model effectiveness in handling both exact and approximate answer matching in Hindi and Marathi QA tasks.
@@ 4.1 Dataset
We utilized the IndicQA dataset, which was constructed by aggregating multiple open-source datasets to develop a SQuAD-style resource for Hindi and Marathi. This dataset includes a significant component of approximately 28,000 samples in these languages, making it one of the largest publicly available datasets in Hindi and Marathi  (Sabane et al., 2024).The datasets are summarized in  Table 2.
@@
Train Set (#sent) Test Set (#sent)
@@ 4.2 Implementation Details
For Mamba and Mamba-2, the fine-tuning configuration included a learning rate of $2\times 10^{-4}$, a batch size of 4 (with gradient accumulation for an effective batch size of 32), 3 epochs, 100 warmup steps, and mixed precision (FP16). The Falcon Mamba model was fine-tuned with start and end predictors, cross-entropy loss, and the Adam optimizer, using a learning rate of $1\times 10^{-4}$, a batch size of 4, and 10 epochs. The Jamba model employed a LoRA configuration with an increased rank of 16 to enhance capacity and processed sequences with a context window of up to 256K tokens. The Zamba model was optimized for long sequences, using a maximum sequence length of 4096 tokens. The Samba model focused on computational efficiency, handling sequences up to 256K tokens with a batch size of 8 (effective batch size of 64 with gradient accumulation), excelling in modeling complex conversational patterns. Finally, the Hymba model incorporated learnable meta tokens and partial sliding window attention, trained with a batch size of 4 and a learning rate of $3\times 10^{-4}$. Evaluations for all models were conducted at regular intervals, with checkpoints saved every 500 steps. At the end of training, the best-performing model was selected. The validation graphs for all models are provided in  Figure 4.
@@ 4.3 Results
This section evaluates model performance on Hindi and Marathi datasets, comparing results before and after fine-tuning. Additionally,  Figure 8 includes sample model responses to highlight qualitative differences in predictions.
4.3.1 Results without fine-tuning
@@
4.3.2 Results With Fine-Tuning
Fine-tuning significantly enhanced model performance across all evaluation metrics, demonstrating the importance of task-specific adaptation for Indic languages. The most improvements were observed in Mamba-2, which exhibited superior alignment with reference answers, improved span localization, and increased robustness in handling complex linguistic structures. These gains highlight the ability of fine-tuned models to better capture long-range dependencies and semantic nuances in Hindi and Marathi question-answering tasks.One key observation is that fine-tuning helped narrow the performance gap between Hindi and Marathi, particularly in models that struggled before adaptation. While Marathi remained more challenging, models showed increased accuracy in capturing token overlaps and generating more contextually relevant responses. Additionally, models that initially performed poorly, such as Falcon Mamba and Hymba, saw only marginal improvements, suggesting fundamental architectural limitations that fine-tuning alone could not overcome. Beyond numerical improvements, fine-tuning also contributed to better lexical and semantic alignment. Models demonstrated enhanced fluency in response generation, with improvements in phrase structure and word selection, reducing inconsistencies seen in non-fine-tuned outputs. However, challenges persisted for certain models, particularly in maintaining coherence when processing long-form text in Marathi. These findings emphasize the necessity of fine-tuning pre-trained models on domain-specific datasets to optimize their performance for low-resource languages. Models results after fine-tuning is provided in  Table 4 , illustrating the extent of improvement across different architectures.
@@ 5 Conclusion
Impact Statement
This research addresses key challenges in developing efficient and context-aware QA systems for Indic languages by leveraging SSMs. It introduces a scalable framework tailored to the complex morphology, syntax, and contextual nuances of these languages. The study advances NLP for low-resource languages, setting a benchmark for multilingual scenarios through SSM-based architectures. By focusing on widely spoken languages like Hindi and Marathi, it enhances accessibility and inclusivity, enabling AI systems to serve underrepresented linguistic communities. The research has real-world applications across sectors such as education, healthcare, governance, and customer support, improving service delivery through localized QA systems. Its emphasis on computational efficiency ensures that these models can be deployed in resource-constrained environments, extending their impact to underserved regions. This work lays a strong foundation for future Indic language QA research and provides a scalable blueprint for adapting similar approaches to other low-resource and linguistically diverse languages worldwide.
@@ Appendix A Dataset Statistics
The scatter plot in  Figure 5 illustrates the relationship between question length and answer length across the Hindi (blue) and Marathi (red) datasets. The visualization highlights a clear positive correlation between the lengths of questions and their corresponding answers, with longer questions often leading to longer answers. Notably, the Marathi dataset shows slightly shorter average lengths compared to the Hindi dataset, reflecting language-specific characteristics. This analysis provides insights into the structural patterns within the dataset, aiding in understanding the data distribution and informing preprocessing decisions for the SSM-based QA model.
Figure 5: Question Length vs Answer Length. This plot compares the lengths of questions and answers in the Hindi (blue) and Marathi (red) datasets, highlighting a correlation between them.”
The scatter plot in  Figure 6 illustrates the relationship between the starting position of answers within the context and the total context length for the Hindi (blue) and Marathi (red) datasets. The visualization shows that as the context length increases, the starting positions of answers become more varied, indicating that longer contexts are more likely to contain answers at different positions. This trend is consistent across both datasets, although the Marathi dataset tends to have shorter contexts on average compared to Hindi. This analysis highlights the diversity in answer positions relative to context length, which is crucial for designing models capable of handling varied input structures in question-answering tasks.
@@ Appendix B Results
The results, as shown in  Figure 9, illustrate the performance metrics for the Hindi and Marathi datasets, comparing outcomes before and after fine-tuning the model. Metrics such as BLEU, ROUGE, F1, Exact Match, and BERT scores are presented for both languages, highlighting the impact of fine-tuning on model performance. The left panel represent the scores without fine-tuning, while the right panels show the improved results after fine-tuning. Across both datasets, fine-tuning consistently improves all evaluation metrics, with notable gains in Exact Match and F1 scores, indicating better alignment between predicted and ground-truth answers. The BERT scores also demonstrate significant improvements, reflecting enhanced semantic understanding. These results emphasize the importance of fine-tuning for adapting the model to the nuances of Hindi and Marathi languages in the QA task.
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Figure 9: Comparison of BLEU, ROUGE, F1, Exact Match, and BERT scores for Hindi and Marathi datasets before and after fine-tuning.

### v2

@@ 3.4 Prompting Strategies for Indic Language based SSMs
SSMs like Mamba and Mamba-2 require careful prompt engineering to fully leverage their question-answering capabilities, especially in low-resource languages such as Hindi and Marathi. One key challenge arises from the linguistic properties of these languages—particularly the Devanagari script—where lexical units often vary in length, leading to non-uniform tokenization. Our approach addresses these challenges by combining the architectural strengths of SSMs with tailored linguistic strategies designed specifically for Devanagari-based languages.
@@
For few-shot prompting Ye and Durrett (2022), The model maintains separate state trajectories for example patterns through its recurrent SSM block.
Figure 3: Examples of structured question-answer pairs in Marathi used for training or evaluating the QA system.
We provide examples of prompts with both positive and negative samples in Appendix  fig. 8, showcasing their use in few-shot prompting, while examples for one-shot prompting are illustrated in  fig. 3. Our Question-Answering pipeline is based on the few-shot prompting framework  Reynolds and McDonell (2021). This approach demonstrates how leveraging paradigms from different prompting techniques leads to improved results during fine-tuning, as evidenced in the subsequent  section 4.
We tend to select the best answer in P where $P\in K;K\geq 0$, are the number of samples generated for that query. These queries are converted into positive and negative samples. These selection of positive and negative samples are done using infomax algorithm Veličković et al. (2018) which introduces stochasticity at the sampling stage but results in more consistent and deterministic outputs by selecting the best among multiple candidates.
@@ 3.5 Evaluation Metrics
To comprehensively evaluate model performance in question-answering tasks, we employed multiple metrics capturing different aspects of accuracy and relevance. EM provides a strict precision measure by evaluating the percentage of predictions that exactly match the ground truth, making it ideal for extractive QA tasks Maalouly (2022). F1 Score balances precision and recall, rewarding partial matches to better reflect token-level similarity, which is especially useful for paraphrased responses Powers (2020). To assess deeper semantic relationships, BERTScore Zhang et al. (2020) leverages contextual embeddings from a pre-trained model, capturing meaning beyond exact word matches. ROUGE evaluates n-gram recall and overlap, ensuring that key phrases from the reference answer are retained even if the phrasing differs Lin (2004). Lastly, BLEU quantifies n-gram precision while incorporating a brevity penalty to prevent overly short predictions, making it particularly effective for assessing fluency and word-level accuracy in extractive QA settings Papineni et al. (2002). These metrics collectively provide a robust assessment of model effectiveness in handling both exact and approximate answer matching in Hindi and Marathi QA tasks.
@@ 4.1 Dataset
We utilized the IndicQA dataset, which was constructed by aggregating multiple open-source datasets to develop a SQuAD-style resource for Hindi and Marathi. This dataset includes a significant component of approximately 28,000 samples in these languages, making it one of the largest publicly available datasets in Hindi and Marathi  Sabane et al. (2024).The datasets are summarized in  table 2.
@@
Train Set (#samples) Test Set (#samples)
@@ 4.2 Implementation Details
For Mamba and Mamba-2, the fine-tuning configuration included a learning rate of $2\times 10^{-4}$, a batch size of 4 (with gradient accumulation for an effective batch size of 32), 3 epochs, 100 warmup steps, and mixed precision (FP16). The Falcon Mamba model was fine-tuned with start and end predictors, cross-entropy loss, and the Adam optimizer, using a learning rate of $1\times 10^{-4}$, a batch size of 4, and 10 epochs. The Jamba model employed a LoRA configuration with an increased rank of 16 to enhance capacity and processed sequences with a context window of up to 256K tokens. The Zamba model was optimized for long sequences, using a maximum sequence length of 4096 tokens. The Samba model focused on computational efficiency, handling sequences up to 256K tokens with a batch size of 8 (effective batch size of 64 with gradient accumulation), excelling in modeling complex conversational patterns. Finally, the Hymba model incorporated learnable meta tokens and partial sliding window attention, trained with a batch size of 4 and a learning rate of $3\times 10^{-4}$. Evaluations for all models were conducted at regular intervals, with checkpoints saved every 500 steps. At the end of training, the best-performing model was selected. The validation graphs for all models are provided in  fig. 4.
@@ 4.3 Results
This section evaluates model performance on Hindi and Marathi datasets, comparing results before and after fine-tuning. Additionally,  fig. 8 includes sample model responses to highlight qualitative differences in predictions.
4.3.1 Pre-Fine-Tuning Performance: Initial Challenge
@@
4.3.2 Post-Fine-Tuning Performance: Improvements & Insights
Fine-tuning significantly enhanced model performance across all evaluation metrics, demonstrating the importance of task-specific adaptation for Indic languages. The most improvements were observed in Mamba-2, which exhibited superior alignment with reference answers, improved span localization, and increased robustness in handling complex linguistic structures. These gains highlight the ability of fine-tuned models to better capture long-range dependencies and semantic nuances in Hindi and Marathi question-answering tasks. One key observation is that fine-tuning helped narrow the performance gap between Hindi and Marathi, particularly in models that struggled before adaptation. While Marathi remained more challenging, models showed increased accuracy in capturing token overlaps and generating more contextually relevant responses.
Additionally, models that initially performed poorly, such as Falcon Mamba and Hymba, saw only marginal improvements, suggesting fundamental architectural limitations that fine-tuning alone could not overcome. Beyond numerical improvements, fine-tuning also contributed to better lexical and semantic alignment. Models demonstrated enhanced fluency in response generation, with improvements in phrase structure and word selection, reducing inconsistencies seen in non-fine-tuned outputs. However, challenges persisted for certain models, particularly in maintaining coherence when processing long-form text in Marathi. These findings emphasize the necessity of fine-tuning pre-trained models on domain-specific datasets to optimize their performance for low-resource languages. Models results after fine-tuning is provided in  table 4 , illustrating the extent of improvement across different architectures.
@@ 5 Conclusion
6 Limitations
Our study presents promising results in applying SSMs for structured question answering in Indic languages; however, there are several limitations to consider. The availability of high-quality, large-scale annotated question-answering datasets for Indic languages remains scarce, affecting model generalizability and hindering further improvements. Although Hindi and Marathi exhibit strong performance improvements post-fine-tuning, other Indic languages with fewer training samples, such as Assamese or Odia, remain underrepresented, limiting the applicability of our models across the full spectrum of Indic languages. Computational efficiency is another concern, particularly for real-time inference on resource-constrained devices. While SSMs offer advantages in handling long sequences, the fine-tuning of large models like Mamba-2 requires significant GPU resources, making widespread deployment challenging. The tokenization process for Indic scripts, especially for languages with complex lexical structures, can introduce alignment errors between context, question, and answer spans, impacting the model’s ability to accurately predict answer positions.
The current models also struggle with answering questions requiring multi-sentence responses or dealing with ambiguity in the provided context, indicating the need for further advancements in contextual reasoning mechanisms. Furthermore, biases in training data may lead the model to favor certain syntactic structures or frequently occurring answer patterns, potentially limiting generalizability to out-of-distribution samples. Our approach also relies heavily on pre-trained SSMs, which may inherit biases from their pre-training corpora. The effectiveness of these models for Indic languages depends on the quality and diversity of the pre-training data, which may not always align with the linguistic characteristics of Indic scripts. Addressing these limitations in future work will involve curating larger and more diverse datasets, optimizing models for lower-resource settings, improving tokenization methods, and developing architectures that enhance contextual reasoning across Indic languages.
@@ Appendix A Dataset Statistics
The scatter plot in  fig. 5 illustrates the relationship between question length and answer length across the Hindi (blue) and Marathi (red) datasets. The visualization highlights a clear positive correlation between the lengths of questions and their corresponding answers, with longer questions often leading to longer answers. Notably, the Marathi dataset shows slightly shorter average lengths compared to the Hindi dataset, reflecting language-specific characteristics. This analysis provides insights into the structural patterns within the dataset, aiding in understanding the data distribution and informing preprocessing decisions for the SSM-based QA model.
Figure 5: Question Length vs Answer Length. This plot compares the lengths of questions and answers in the Hindi (blue) and Marathi (red) datasets, highlighting a correlation between them."
The scatter plot in  fig. 6 illustrates the relationship between the starting position of answers within the context and the total context length for the Hindi (blue) and Marathi (red) datasets. The visualization shows that as the context length increases, the starting positions of answers become more varied, indicating that longer contexts are more likely to contain answers at different positions. This trend is consistent across both datasets, although the Marathi dataset tends to have shorter contexts on average compared to Hindi. This analysis highlights the diversity in answer positions relative to context length, which is crucial for designing models capable of handling varied input structures in question-answering tasks.
@@ Appendix B Results
The results, as shown in  fig. 9, illustrate the performance metrics for the Hindi and Marathi datasets, comparing outcomes before and after fine-tuning the model. Metrics such as BLEU, ROUGE, F1, Exact Match, and BERT scores are presented for both languages, highlighting the impact of fine-tuning on model performance. The left panel represent the scores without fine-tuning, while the right panels show the improved results after fine-tuning. Across both datasets, fine-tuning consistently improves all evaluation metrics, with notable gains in Exact Match and F1 scores, indicating better alignment between predicted and ground-truth answers. The BERT scores also demonstrate significant improvements, reflecting enhanced semantic understanding. These results emphasize the importance of fine-tuning for adapting the model to the nuances of Hindi and Marathi languages in the QA task.
@@
Figure 9: Comparison of BLEU, ROUGE, F1, Exact Match, and BERT scores for Hindi and Marathi datasets before and after fine-tuning.